Kanye wants to save ​$14,000 in 7 years by making monthly payments into an ordinary annuity for a down payment on a condominium at the shore. If the annuity pay

0.6​% monthly​ interest, what will his monthly payment​ be?

Round up to the nearest cent
Answer #1

Future value = pmtx ((1+r)^n-1))/ PMT = Save each month = PMT= FV + ((1+r)^n-1))/r $128.67 [ 14,000/(((1.006484)-1)/0.006)]